# Critical dynamics and its interferometry in the one-dimensional p-wave-paired Aubry-André-Harper model

### v1

@@ IV.1.1 Defect behavior after one-way quenching
Additional attention is placed on the fast quench dynamics. In Fig. 6 (a), we collect the value of plateau $n_{\text{su}}$ as a function of the reverse of initial potential strength within very fast quench rate. A fitting for $1/2-n_{\text{su}}=\alpha/V_{i}$ yields the fitting parameter $\alpha\approx 0.425\pm 0.0004$. Furthermore, in Fig. 6(b), we show $(n_{\text{su}}-n_{\text{ex}})V_{i}$ as a function of $V_{i}^{2}\tau_{Q}$ by varying initial potential strength $V_{i}$. In the SR, $(n_{\text{su}}-n_{\text{ex}})V_{i}$ with different $V_{i}$ share squared behavior of the scaled quench time $V_{i}^{2}\tau_{Q}$, which corresponds to with the fitting parameter $C\approx 0.10\pm 0.005$. In PSR, the turning point of SR and PSR, being proportional to the reciprocal of the squared initial potential strength, $\tau_{Q}^{S}\sim V_{i}^{-2}$, is reflected as an obvious change of the behavior of combined defects’ density, $(n_{\text{su}}-n_{\text{ex}})V_{i}$, in Fig. 6(b). The scaling behavior of defect density is similar to that in the transverse Ising model, which corresponds to the similar critical behavior with $\nu,z=1$.
@@
IV.2 Round-trip quench and the necking phenomena
@@
Figure 8: The period of oscillatory defects obtained by varying to-and-fro quench time ratio $R$ and turning point $V_{rt}$. (a) $1/T_{Q}$ versus $1+R$ with $V_{rt}=1$. A linear model fitting of $1/T_{Q}=a(1+R)$ yields $a\approx 0.524\pm 0.0005$. (b) $1/T_{Q}$ versus $\left(V_{rt}-C\right)^{2}$ with $R=1$. A linear model fitting of $1/T_{Q}=b\left(V_{rt}-C\right)^{2}$ yields $b\approx 0.347\pm 0.004$ and $C\approx 1.742\pm 0.009$.
Figure 9: (a) the density of defect as a function of the quench time $\tau_{Q}$ by setting the ratio of to-and-fro quench time $R=1$. The narrow neck is exposed by curves with different turning points $V_{rt}$. (b) The reciprocal of neck point $1/\tau_{Q}^{\text{neck}}$ versus $1+R$ with turning point $V_{rt}=1$. A linear fitting of $1/\tau_{Q}^{\text{neck}}=\lambda(1+R)$ yields $\lambda\approx 0.05\pm 0.0001$.
@@ IV.2.1 Interference effect at $V_{rt}=0$ and $R=1$
The final defect’s density as a function of the quench time $\tau_{Q}$ is shown in Fig. 7. After the round-trip protocol, the defect’s density exhibits oscillatory behavior with a steady period $T_{Q}\approx 0.943$ because of the dynamical phase. Interestingly, a narrow neck appears in the oscillatory defect density after the round-trip protocol. To identify the position of the narrow neck, the amplitude $M$ is introduced by the differences between two interpolation functions from wave peaks and troughs(Fig. 7 (b)). The neck position $\tau_{Q}^{\text{neck}}$, defined as the minimal of the amplitude $M$, is well located between KZ-I and KZ-II. In this case, we suggest the narrow neck is a phenomenon ascribed as the changing dynamical exponent (9). Notably, the interference effect is hard to observe in the SR and PSR because the period $T_{Q}$ is larger than the scale of those regimes.
@@ IV.2.2 Interference effect for varying $V_{rt}$ and $R$
Different turning points $V_{rt}$ and the quench times ratio $R$ are investigated for further observing the interference effect. The oscillation periods, as functions of the to-and-fro quench times ratio $R$ and the turning point $V_{rt}$, are collected in Fig. 8. In Fig. 8 (a), the turning point $V_{rt}$ is set to 00, and the final defects’ density is obtained by varying $R$. A linear fitting of the inverse of period $1/T_{Q}$ as a function of $1+R$ gives the proportionality coefficient $\alpha\approx 0.524\pm 0.0005$. In Fig. 8 (b), with a fixed to-and-fro quench times ratio $R=1$, a linear fitting of $1/T_{Q}$ as a function of $(V_{rt}-C)^{2}$ yields the proportionality coefficient $\beta\approx 0.347\pm 0.004$ and the constant $C\approx 1.742\pm 0.009$. These numerical results determine a relationship for the oscillation period:
$T_{Q}\approx\frac{0.695}{(1+R)(V_{rt}-C)^{2}},$ (22)
which is similar to the oscillation period relationship of defects in the transverse field Ising model [10].
Furthermore, the neck position is also investigated by varying $V_{rt}$ and $R$. In Fig. 9 (a), setting the quench time ratio at $R=1$, the excitation probability as a function of $\tau_{Q}$ exhibits a common narrow neck with different turning points $V_{rt}$. The stationary neck position under changing $V_{rt}$ reveals that it is irrelevant to the dynamical phase. In Fig. 9 (b), we set the first linear ramps’ quench time $\tau_{Q}$ and adjust the second one $\tau_{Q}^{{}^{\prime}}$ to control the parameter $R$. By varying the ratio of to-and-fro quench times $R$, the neck position behaves as a function of the reciprocal of $1+R$, $\tau_{Q}^{\text{neck}}\sim(1+R)^{-1}$. A linear model fitting of $1/\tau_{Q}^{\text{neck}}$ as a function of $1+R$ yields a proportionality coefficient $\lambda^{-1}\approx 0.05\pm 0.0001$. Setting $R=1$, corresponding to the same to-and-fro quench rate, the neck position is observed at the turning point between KZ-I and KZ-II. By increasing the returning quench rate, corresponding to a larger $R$, the system enters KZ-II earlier during the return process, causing the neck position to appear sooner than at $R=1$, and vice versa.
In conclusion, the changing dynamical exponent (9) is disclosed by a narrow neck in the oscillatory defect density after the round-trip quench protocol. Additionally, the position of narrow neck is well located between two KZ sub-regimes, which provides us a way to find the shifting point of two KZ power-law scaling.
@@ IV.3 Varying the model parameter $\Delta$
To show the universality of the alternation of the KZ exponent behavior in the one-way quench protocol and the neck phenomenon in the round-trip quench protocol. It is essential to investigate different p-wave superconductivity pairing amplitude $\Delta$.
The final density of defects is collected as a function of quench time after one-way and round-trip quench protocols in the system with different the p-wave pairing amplitude $\Delta$. One should notice that the system across the extended-critical transition point will not get excitation because the energy gap is open at this point $V_{c_{2}}=\left|J-\Delta\right|$.
@@
In Fig.10 (a), the neck positions, extracted from the round-trip quench, as a function of p-wave superconductivity pairing amplitude $\Delta$ show a power-law behavior $\tau_{Q}^{\text{neck}}\sim\Delta^{-h}$ with $h=1.39\pm 0.03$. In Fig.10 (b), we obtain the final excitation density populated after one-way quenching versus the quench time $\tau_{Q}$ by varying p-wave superconductive pairing amplitudes $\Delta$. The fits of $n_{\text{ex}}\sim\tau_{Q}^{-r}$ yields $r\approx 0.5$ for each $\Delta$ for relevant fast quench rate, corresponding to the dash lines. The fitting is only applied for KZ-I because the power-law scaling of KZ-II occurs when there is a long quenching period. For different $\Delta$, neck positions are observed at the turning point between two sub-regimes exhibiting different power-law scaling behavior. By multiplying the scaled factor $\Delta^{1.39}$ by the quench time, different power-law scaling is collapsed to a universal curve in Fig.10 (c). A common neck position can divide the KZ regime into two sub-regimes, KZ-I and KZ-II, which show the universality of two sub-regimes phenomena in the Aubry-André-Harper(AAH) model.
@@ V Conclusion
In the gapless localized phase of the p-wave paired Aubry-André-Harper model, we have found that the dynamical exponent is continuously dependent on the potential strength. As the main feature, two plateaus have been identified in the dynamical exponent profile: one in the vicinity of the quantum phase transition point and another away from it. This two-plateau structure of the dynamical exponent turns out to influence the critical dynamics deeply. In the one-way quench protocol, we illustrate that the usual KZ regime consists of two sub-regimes with different KZ exponents as a consequence. While in the round-trip quench protocol, we demonstrate a narrow neck in the oscillatory density of defects as an interference effect of two critical dynamics. By adjusting the to-and-fro quench time ratio and model parameters, we show how the position of the neck can be used to quantitatively determine the turning point of the two KZ sub-regimes.

### v2

@@ IV.1.2 Final defects with fast quench rate
Additional attention is placed on the fast quench dynamics. In Fig. 6(a), we collect the value of plateau $n_{\text{su}}$ as a function of the reverse of initial potential strength within very fast quench rate. A fitting for $1/2-n_{\text{su}}=\alpha/V_{i}$ yields the fitting parameter $\alpha\approx 0.425\pm 0.0004$. Furthermore, in Fig. 6(b), we show $(n_{\text{su}}-n_{\text{ex}})V_{i}$ as a function of $V_{i}^{2}\tau_{Q}$ by varying initial potential strength $V_{i}$. The curves with different $V_{i}$ share a universal scaling $(n_{\text{su}}-n_{\text{ex}})V_{i}\sim V_{i}^{2}\tau_{Q}$, which corresponds to Eq. (15) with the fitting parameter $K\approx 0.10\pm 0.005$. In PSR, the turning point of SR and PSR, being proportional to the reciprocal of the squared initial potential strength, $\tau_{Q}^{S}\sim V_{i}^{-2}$, is reflected as an obvious change of the behavior of combined defects’ density, $(n_{\text{su}}-n_{\text{ex}})V_{i}$, in Fig. 6(b). The scaling behavior of defect density is similar to that in the transverse Ising model, which corresponds to the similar critical behavior with $\nu,z=1$.
@@
IV.2 Round-trip quench protocol
@@
Figure 8: The period of oscillatory defects obtained by varying to-and-fro quench time ratio $R$ and turning point $V_{rt}$. (a) $1/T_{Q}$ versus $1+R$ with $V_{rt}=1$. A linear model fitting of $1/T_{Q}=a(1+R)$ yields $a\approx 0.524\pm 0.0005$. (b) $1/T_{Q}$ versus $\left(V_{rt}-C\right)^{2}$ with $R=1$. A fitting of $1/T_{Q}=b\left(V_{rt}-C\right)^{2}$ yields $b\approx 0.347\pm 0.004$ and $C\approx 1.742\pm 0.009$.
Figure 9: (a) the density of defect as a function of the quench time $\tau_{Q}$ by setting the ratio of to-and-fro quench time $R=1$. The narrow neck is exposed by curves with different turning points $V_{rt}$. (b) The reciprocal of neck point $1/\tau_{Q}^{\text{neck}}$ versus $1+R$ with turning point $V_{rt}=0$. A linear fitting of $1/\tau_{Q}^{\text{neck}}=\lambda(1+R)$ yields $\lambda\approx 0.05\pm 0.0001$.
@@ IV.2.1 Interference effect at $V_{rt}=0$ and $R=1$
The final defect’s density as a function of the quench time $\tau_{Q}$ is shown in Fig. 7. After the round-trip protocol, the defect’s density exhibits oscillatory behavior with a steady period $T_{Q}\approx 0.943$ because of the dynamical phase. Interestingly, a narrow neck appears in the oscillatory defect density after the round-trip protocol. To identify the position of the narrow neck, the amplitude $M$ is introduced by the differences between two interpolation functions from wave peaks and troughs (Fig. 7(b)). The neck position $\tau_{Q}^{\text{neck}}$, defined as the minimum of the amplitude $M$, is well located between KZ-I and KZ-II. In this case, we suggest the narrow neck is a phenomenon ascribed as the changing dynamical exponent (9). Notably, the interference effect is hard to be observed in the SR and PSR because the period $T_{Q}$ is larger than the scale of those regimes.
@@ IV.2.2 Interference effect for varying $V_{rt}$ and $R$
Different turning points $V_{rt}$ and the quench times ratio $R$ are investigated for further investigating the interference effect. The oscillation periods, as functions of the to-and-fro quench times ratio $R$ and the turning point $V_{rt}$, are collected in Fig. 8. In Fig. 8(a), the turning point $V_{rt}$ is set to 00, and the final defects’ density is obtained by varying $R$. A linear fitting of $1/T_{Q}=a(1+R)$ yields the proportionality coefficient $a\approx 0.524\pm 0.0005$. In Fig. 8 (b), with a fixed to-and-fro quench times ratio $R=1$, a fitting of $1/T_{Q}=b(V_{rt}-C)^{2}$ provides the proportionality coefficient $b\approx 0.347\pm 0.004$ and the constant $C\approx 1.742\pm 0.009$. These numerical results determine a relationship for the oscillation period:
$T_{Q}\approx\frac{5.78\pm 0.09}{(1+R)(V_{rt}-C)^{2}},$ (22)
which is similar to the results in the transverse field Ising model [14].
Furthermore, the neck position is also investigated by varying $V_{rt}$ and $R$. In Fig. 9(a), setting the quench time ratio at $R=1$, a distinct narrow neck is disclosed by varying the turning point $V_{rt}$. The position of this neck position is independent of $V_{rt}$, which reveals that the neck is irrelevant to the dynamical phase. In Fig. 9(b), we set the first linear ramps’ quench time $\tau_{Q}$ and adjust the second one $\tau_{Q}^{{}^{\prime}}$ to control the to-and-fro quench times ratio $R$, with $V_{rt}=0$. The neck position is found to be a function of the reciprocal of $1+R$, i.e. $\tau_{Q}^{\text{neck}}\sim(1+R)^{-1}$. The neck position with $R=1$ corresponds to the turning point between KZ-I and KZ-II (Fig. 7(b)). An increased returning quench rate (associated with larger $R$) induces earlier entry into the KZ-II regime during the return process. This advancement shifts the neck position to occur at an earlier stage compared to the $R=1$ case, with reduced $R$ conversely delaying this event.
In conclusion, the two-plateau structure of the dynamical exponent (9) is reflected by a narrow neck in the oscillatory defect density after the round-trip quench protocol. Additionally, the position of narrow neck is well located between two KZ sub-regimes, which can be utilized to determine the turning point of KZ-I and KZ-II.
@@ IV.3 Varying the model parameter $\Delta$
Then we investigate the influence of the two-plateau structure (9) by adjusting the system parameters $\Delta$. The final density of defects is collected as a function of quench time after one-way and round-trip quench protocols in the system with different the p-wave pairing amplitude $\Delta$. One should notice that the system across the extended-critical transition point will not get excitation because the energy gap is open at this transition point $V_{c_{2}}=\left|J-\Delta\right|$.
@@
In Fig. 10(a), the neck positions, extracted from the round-trip quench, as a function of p-wave superconductivity pairing amplitude $\Delta$ show a power-law behavior $\tau_{Q}^{\text{neck}}\sim\Delta^{-h}$ with $h=1.39\pm 0.03$. In Fig. 10(b), we obtain the final excitation density populated after one-way quenching versus the quench time $\tau_{Q}$ by varying p-wave superconductive pairing amplitudes $\Delta$. The fits of $n_{\text{ex}}\sim\tau_{Q}^{-r}$ yields $r\approx 0.5$ for each $\Delta$ within moderately fast quench rate, corresponding to the dash lines. The fitting is only applied for KZ-I because the power-law scaling of KZ-II occurs when there is a long quenching period. For different $\Delta$, neck positions are observed between the turning point of two sub-regimes with different KZ scaling behavior. By multiplying the scaled factor $\Delta^{1.39}$ by the quench time, different power-law scaling is collapsed to a universal curve in Fig. 10(c). The common neck position can divide the KZ regime into two sub-regimes, which shows that the influence the critical dynamics holds universally across all model parameters $\Delta$.
@@ V Conclusion
In the gapless localized phase of the p-wave paired Aubry-André-Harper model, we have found that the dynamical exponent is continuously dependent on the potential strength. As the main feature, two plateaus have been identified in the dynamical exponent profile: one in the vicinity of the quantum phase transition point and another away from it. This two-plateau structure of the dynamical exponent turns out to influence the critical dynamics deeply. In the one-way quench protocol, we illustrate that the usual KZ regime consists of two sub-regimes with different KZ exponents as a consequence. While in the round-trip quench protocol, we demonstrate a narrow neck occurring in the oscillatory density of defects as an interference effect of two critical dynamics. The phenomenon of narrow neck is a direct reflection of the two-plateau structure of the dynamical exponent, which means such an interferometry provides a valuable method for probing the critical behavior of the quasiperiodic system. By adjusting the to-and-fro quench time ratio and model parameters, we show how the position of the neck can be used to quantitatively determine the turning point of the two KZ sub-regimes.
VI ACKNOWLEDGMENTS
We thank Jian-Song Pan, Yan He, B. Hetényi, L. Sanchez-Palencia, G. Roosz, and Uwe R. Fischer for insightful discussions. This work is supported by NSFC under Grant No. 11074177.